OpenStudy (zion):

if g(x)=log base7 X, then g^-1(x)=?

OpenStudy (anonymous):

g^-1(x)=7^x

OpenStudy (zion):

how?

OpenStudy (anonymous):

nah, its \[\log_{x} 7\]

OpenStudy (anonymous):

working y=logbase7 X 7^y=x f^-1(x)=7^y

OpenStudy (anonymous):

youre asking about \[1 \div g (x) \] right?

OpenStudy (zion):

hashir got the correct answer.

OpenStudy (anonymous):

yeah i am always right!!!!!!

OpenStudy (zion):

hahaha. stay humble dude! hahahahahaha!

OpenStudy (anonymous):

hahahahahah jk fb id plz :D

OpenStudy (zion):

www.facebook.com/paulguarnes2 ... you cant add me on my main account coz its already full.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

anyway, take a look at this:: http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

it is inverse boy !!!!!!!!!

OpenStudy (anonymous):

i think u need to learn functions

OpenStudy (anonymous):

+ log too

OpenStudy (anonymous):

right. got lots to learn. about understanding the coded lingo here. :P

OpenStudy (zion):

The questions that im asking are from my reviewer, and they got the correct answer, but the solutions are not provided. Hashir got the correct answer, all i wanna learn is how to get the answer.

OpenStudy (anonymous):

Yeah I got the Same

OpenStudy (anonymous):

Ok let me re-type the whole thing

OpenStudy (anonymous):

\[g(x) = log_7 x\] Now Let g(x) = y \[y = log_7x\]

OpenStudy (anonymous):

Hence to get Inverse we need to turn x as a function of y

OpenStudy (anonymous):

\[x = 7^y\]From\[\large{log_{base}x =y => x = {base}^y}\]

OpenStudy (zion):

the answer is 7^x

OpenStudy (anonymous):

Now we have the Setup that makes \[g^{-1}(x) = 7^x\]is the Inverse

OpenStudy (anonymous):

yeah you got it : )